Annual savings due to an energy efficiency project have a most likely value of $30,000. The high


estimate of $40,000 has a probability of 0.2, and the low estimate of $20,000 has a probability of 0.30.


What is the expected value for the annual savings?

1
Expert's answer
2021-12-03T15:18:02-0500

The high estimate of $40,000 has a probability of 0.25, and the low estimate of $20,000 has a probability of .35. Thus, the estimate of $30,000 has the probability of (1-0.25-0.35)=0.40.

Now, the expected value for the annual savings is the weighted average of the annual savings.

Hence, 

the expected value for the annual savings = 0.40×30000+0.25×40000+0.35×20000 = 29000


Therefore, the expected value for the annual savings is $29000.